angle a and angle b are complementary angles.if one angle is 15 degree more than 2 times the other.find the measure of two angles
Answer by PRMath(133)

First you have to know two definitions. Well, you only have to know one for this problem, but you should probably learn the 2nd just to be thorough.
Definition 1: Complementary angles are two angles whose sum is 90 degrees.
Definition 2: Supplementary angles are two angles whose sum is 180 degrees.

For this problem, we'll work with the definition that says two complementary angles have a sum of 90 degrees.

Soooo, here are the facts from your problem: if one angle is 15 degree more than 2 times the other.find the measure of two angles.

Let's let the larger angle equal this: 15 + 2(x) (<--See how it is 15 degrees MORE than 2 times the other?)
Let's let the smaller angle equal: x


SO now our total equation is:

15 + 2(x) + x = 90
3x + 15 = 90 (combined like terms)
3x = 75 (subtracted 15 from both sides)
x = 25 (divided both sides by 3)

Now we know that one angle is 25. The other angle must add to 25 to make 90 degrees, so 90 - 25 = 65.

Therefore, your two angles are 25 and 65 degrees.

Does this check out? Let's see...
First: 25 + 65 = 90 Therefore, this checks out.

Second: The angle that is 65 degrees must be 15 degrees more than twice the other. So, let's take twice the other...... 25 * 2 = 50. And, let's add 15....50 + 15 = 65. Therefore YES, the 2nd angle is 15 more than 2 times the angle that was 25 degrees.

I hope this is helpful. :-)